OpenStudy (anonymous):

While skiing, Ellen encounters a ledge that sends her flying horizontally at a rate of 12 m/s. How far away will she land if the ledge is 7 meters high?

OpenStudy (anonymous):

12 x sqrt(1.4)

OpenStudy (anonymous):

Could you explain a little bit because Im not sure I understand

OpenStudy (anonymous):

7 meters high will mean that she will fall 7 meters down. How much time does this take? s = 1/2 a t^2 7 = 1/2*9.81*t^2 t = 1.19 horizontal distance = v*t = 12*1.19 = 14.33 meters

OpenStudy (anonymous):

wouldnt t= 1.42?

OpenStudy (anonymous):

no, because it's t^2 = 1.42

OpenStudy (anonymous):

Okay, could you type that out please?

OpenStudy (anonymous):

s = 1/2 a t^2 7 = 1/2*9.81*t^2 t^2 = 1.42 t = 1.19